A mixture of 10ml CH4, C2H4 and C2H2 has a vapour density of 11.3 . Mixture contains x ml of CH4,  y ml of C2H4 and z ml of C2H2. When 30 ml of oxygen are sparked together over aqueous KOH, the volume contracts to 5.5ml and then disappears when pyrogallol is introduced. If volumes are measured in same conditions of pressure, temperature and humidity, value of x, y and z are
 

Grade:11

1 Answers

Vikas TU
3 years ago
Let the volume of CH4 be a ml 
Volume of C2H4 = b ml
Hence volume of CO2 = (10 - a - b) ml
After the explosion, there was contraction of 17ml which means that when the explosion takes place at that time water is in gaseous state and after the explosion water vapors converted into liquid state.Also the volume of liquid water is assumed to be zero.
Hence 2a + 2b = 17    ........(1)
Because KOH absorbs CO2 gas when the gaseous mixture is treated with KOH ,there is a further reduction in volume of about 14 ml, hence volume of CO2 in gaseous mixture is 14 ml.
10 - a - b + a + 2b = 14.. (2)
=> 10 + b = 14
=> b = (14-10)ml = 4 ml
By using this value in equation (1), we get
2a + 2b = 17
=> 2a = 17- 2 x 4 = 17-8 = 9
=>a = 9/2 = 4.5 ml
So, volume of CH4 in the mixture = 4.5 ml
Also,Volume of C2H4 in the mixture = 4 ml
That of CO2 in the mixture= (10 - 4 - 4.5) = 1.5 ml